# NeRF-Supervised Feature Point Detection and Description

### v1

@@ Abstract
Feature point detection and description is the backbone for various computer vision applications, such as Structure-from-Motion, visual SLAM, and visual place recognition. While learning-based methods have surpassed traditional handcrafted techniques, their training often relies on simplistic homography-based simulations of multi-view perspectives, limiting model generalisability. This paper introduces a novel approach leveraging neural radiance fields (NeRFs) for realistic multi-view training data generation. We create a diverse multi-view dataset using NeRFs, consisting of indoor and outdoor scenes. Our proposed methodology adapts state-of-the-art feature detectors and descriptors to train on NeRF-synthesised views supervised by perspective projective geometry. Our experiments demonstrate that the proposed methods achieve competitive or superior performance on standard benchmarks for relative pose estimation, point cloud registration, and homography estimation while requiring significantly less training data compared to existing approaches.
@@ 1 Introduction
Feature point detection and description under different scene viewpoints is a common starting point for many multi-view problems including Structure-from-Motion [1], visual SLAM [2, 3], or visual place recognition [4, 5]. In the past decade, different learning-based approaches to this problem [6, 7, 8, 9] have replaced handcrafted techniques in many applications [10, 11, 12]. Crucially, most of these models can be fine-tuned in a self-supervised manner on any single-view dataset. This is achieved by applying different homography warpings to the training data, simulating different viewpoints of the same scene with known point-to-point "ground-truth" mappings. While this training scheme is simple and flexible, the generated homography warpings are a crude simplification of multi-view perspective, which can lead to limited model generalisability.
This paper aims at leveraging image synthesis with neural radiance fields (NeRFs) as a more realistic way of generating multi-view training data for feature detection and description models (see Fig. 1). Since NeRFs require multi-view data to synthesise novel views, we can no longer rely on the single image datasets typically used for training the above mentioned homography-based methods. We therefore create our own dataset consisting of both indoor and outdoor image sequences around static scenes, and reconstruct all of them with NeRFacto [13]. This enables the generation of arbitrary viewpoints of each scene consistent with a pinhole projective model with known point-to-point mappings via point re-projection. We propose a general methodology to upgrade state-of-the-art homography-based methods to train them on projective views synthesised from NeRF-type algorithms. Our contributions are as follows:
  • We create a new multi-view dataset consisting of images from 10 different indoor and outdoor scenes, and a total of 10000 NeRF-synthesised views from these scenes with corresponding depth maps, intrinsic and extrinsic parameters.
@@
  • We re-train adapted versions of SuperPoint [6] and SiLK [14] using our NeRF-synthesised data and compare them against the original baselines trained on the much larger MS-COCO dataset. We outperform the original baselines for relative pose estimation on ScanNet and YFCC100M datasets, similar performance for pair-wise point cloud registration, while only slightly under-performing on the HPatches homography estimation benchmark.
@@
Figure 1: Visual representation of multi-view data. Learning-based detectors and descriptors achieve supervision on single-view datasets by simulating different viewpoints through homograhpic warpings to the input image $I$ (see Fig. 0(a)) resulting in $I^{{}^{\prime}}$ (see Fig. 0(b)). While we achieve supervision by directly sampling a NeRF rendered image from a different viewpoint (see Fig. 0(c)).
@@ 2 Related Work
Research on multi-view feature detection and description has a large focus on creating representations invariant to geometric transformations (e. g. scale, rotation, affinity) and to illumination conditions, so that the same scene point can be reliably recognised and matched regardless of its viewpoint. Classic methods achieved this by handcrafting feature extractors that incorporate these invariance properties by design. Some of these decade-old methods stood the test of time and are still widely used today [15, 16]. More recently, the success of deep feature extraction enabled neural networks to learn these invariance properties from training data. Large open-source datasets [17, 18, 19] have been extensively employed to train either learning-based interest point detectors and descriptors [6, 20, 14, 21], or detector-free local feature matchers [22, 23, 24].
Some of these methods [6, 14] are trained in a self-supervised manner on uncalibrated RGB single-view datasets such as MS-COCO [17] by simulating multiple views as homography warpings. Other methods [20, 21] are trained with additional supervision using calibrated multi-view RGB-D datasets such as ScanNet [18]. These two classes of methods have pros and cons. The multi-view data generated by self-supervised methods is not representative of all possible viewpoint changes (homographies only model planar scenes or pure rotations) and thus it does not fully generalise to some application scenarios. On the positive side, they can be trained on arbitrarily large number of transformations and are also extremely flexible as they can be fine-tuned on any uncalibrated single-view dataset. Fully supervised methods are trained on more realistic multi-view data that covers full projective view changes with occlusions, however, the required training data is significantly more complex to obtain. The variety of scenes and viewpoint changes in the training data is also limited by the practicalities of real data acquisition. To get the best of both worlds, we propose using NeRF reconstructions obtained from RGB image sequences, which are both simple to acquire and enable the synthetic generation of arbitrary projective viewpoints beyond homographies.
Progress on image synthesis with NeRF and its variants [25, 26, 27] has exploded in the past years, with very fast improvements in computational efficiency, image synthesis quality, reconstruction accuracy, and input data requirements. As these techniques enable synthesising arbitrary viewpoints of a reconstructed scene, they can be used for rich data augmentation [28]. Most related to our work, they have been utilised to self-supervise multi-view methods for learning viewpoint invariant object descriptors [29], stereo disparity [30], and optical flow [31]. However, to the best of our knowledge NeRF self-supervision has been under-explored for point feature detection and description.
@@ 3.1 NeRF Dataset
We build a NeRF dataset for 10 indoor and outdoor scenes, containing in total 10000 synthetic images with their corresponding intrinsic/extrinsic parameters and depth maps. We note that this dataset is extremely small compared to what is typically used to train state-of-the-art feature detection (30 times smaller than MS-COCO, 250 times smaller than ScanNet). Our goal is to show that leveraging NeRF self-supervision enables state-of-the-art performance on multiple benchmarks with significantly less training data.
We captured four indoor and one outdoor scenes with an IPhone 10 at 4K resolution, one indoor and one outdoor scene with a Samsung A52 at 4k resolution. We also utilise two blender generated indoor scenes with images sourced from [32] and one outdoor scene publicly available through [33] as they are scenes that consist of high quality RGB images, while being are rich with features. While different training sets could be constructed using existing publicly available data, we focus on a trade-off between scene diversity and minimal dataset size.
We obtain camera poses for all images using COLMAP [34] and use them to generate synthetic views. We trained a Nerfacto model for each scene, which is part of the NeRFStudio framework [13] and merges improvements from a range of NeRF implementations [35, 33, 36, 37]. For each scene we rendered 1000 synthetic views and corresponding depth maps with $44^{\circ}$ field of view and $640\times 480$ resolution. We use NeRFStudio’s interactive real-time viewer to define custom camera trajectories for each scene that ensure diversity in viewpoints while maintaining global scene overlap. Trajectories consist of uniformly sampled smooth trajectories around the scene coupled with small rotations around the camera centre.
(a) Custom Camera trajectory
(b) Camera 2 Viewpoint
Figure 2: Customised Camera Trajectory. A visual representation of a custom camera trajectory in NeRF Studio’s [13] interactive real-time viewer. Fig. 1(b) demonstrates the viewpoint of Camera 2 in Fig. 1(a).
@@ 3.2 NeRF Point Re-Projection
$\mathbf{p}=\mathcal{K^{\prime}}\frac{\mathbf{P}{{}^{{}^{\prime}}}}{\mathbf{P}{{}^{{}^{\prime}}_{z}}},\quad\mathbf{P}{{}^{{}^{\prime}}}=[R^{\prime T}|-R^{\prime T}\mathbf{t^{\prime}}]\mathbf{P}.$ (2)
Assuming that reconstructed depths $d$ are accurate, the above equations enable self-supervising a point detector, by designing a loss function that promotes a set of re-projected points $\mathbf{p}^{\prime}$ detected in view $I$ to align with points independently detected on view $I^{\prime}$. However, many distinctive interest points lie on scene edges at a depth discontinuity. In this case any small deviation in $\mathbf{p}$ can result in large $d$ variations corresponding to background and foreground depths. To make re-projection stable around edges, we focus on matching foreground objects only (minimising potential view occlusions). This can be achieved in a very simple manner by computing $d$ as the minimum depth within a $5\times 5$ pixel window centred in $\mathbf{p}$ (see example in Fig. 3).
@@ 4.1 SiLK-PrP
In the original SiLK approach, image pairs with ground-truth bi-directional correspondences are generated on-the-fly during training with homography warpings. We rely on our pre-rendered NeRF dataset described in section 3.1, to avoid significant computational overhead caused by on-the-fly image rendering. During training we randomly sample image pairs $I,I^{{}^{\prime}}$ from the same scene, Eqs. 1 and 2 are then utilised during training to generate dense ground-truth point correspondences, discretised at pixel level.
@@ 4.2 SuperPoint-PrP
SuperPoint is a self-supervised method that goes through multiple stages of training. Firstly, an encoder with a detector head is trained in a fully supervised manner to detect corners on a synthetic dataset with simple geometric shapes. The pre-trained detector, called MagicPoint, is used to generate pseudo-ground truth interest points on a real dataset (e.g. MS-COCO) using a process called homographic adaptation. This involves generating several homography warped copies of a training image, passing them through the trained detector, unwarping all predictions back to the original training image, and finally aggregating all unwarped predictions to generate the pseudo label. This process can be repeated several times, where after each round of homographic adaptation, the detector head is further trained, updating the pseudo-ground truth labels, while improving the detector head generalisability. After a final round of homographic adaptation, a descriptor head is added to the model, which is jointly trained with the whole model with an additional descriptor loss term on the real dataset resulting in the SuperPoint model.
With the NeRF training data approach, we use the same architecture of the original SuperPoint model. However, modify the homographic adaptation process as well as the descriptor loss to handle our NeRF training data and Point Re-Projection process.
@@ 4.2.1 Projective Adaptation
Similarly to our SiLK-PrP approach, we consider pre-rendered NeRF images as training data. In this case, we exploit the fact that training sequences for each scene are are generated along a continuous trajectory.
We randomly sample from a window of 20 consecutive rendered NeRF images $I_{i\rightarrow i+19}$. We take pseudo ground-truth points for initial image $I_{i}$ and random sample 15 of the remaining images as the warped versions of $I_{i}$.
The probability heatmap $\mathbf{{H}_{i}}$ for image $I_{i}$ is then computed, followed by randomly extracting 15 images out of the subsequent images. For each of the randomly selected image $I_{r}$, their probability heatmap $\mathbf{{H}_{r}}$ is computed, and interest points $\mathbf{{p}_{r}}$ are extracted from each heatmap, while using non-maximum suppression to filter out feature points that are too close to each other.
Once interest points $\mathbf{{p}_{r}}$ are obtained, the PrP process is applied to project the interest points aligning them with the initial frame ${I}_{i}$, for clarity the projected interest points can be denoted as $\mathbf{{p}^{{}^{\prime}}_{i}}$.
Having obtained the projected interest points $\mathbf{{p}^{{}^{\prime}}_{i}}$, we generate a mask denoted as $\mathbf{{B}_{i^{\prime}}}$, for each interest point in $\mathbf{{p}^{{}^{\prime}}_{i}}$ coordinates on the binary mask, a 3x3 patch centred around the interest point before projection is extracted from the probability heatmap $\mathbf{{H}_{r}}$, the patch is then applied on the binary heatmap $\mathbf{{B}_{i^{\prime}}}$. Thus, mask $\mathbf{{B}_{i^{\prime}}}$, serves as the probability heatmap of rendered image $I_{r}$ projected onto the input render frame ${I}_{i}$.
Finally, the input’s rendered image probability heatmap $\mathbf{{H}_{i}}$ and mask $\mathbf{{B}_{i^{\prime}}}$ for each frame, are aggregated together, non-maximum suppression is applied once more to the final aggregated heatmap to obtain pseudo ground-truth interest points for the input rendered image $I_{i}$.
@@ 4.2.2 Descriptor Loss
SuperPoint’s [6] loss consists of an interest point detection loss $\mathcal{L}_{p}$ and the other for the descriptor $\mathcal{L}_{d}$.
The interest point detection loss is a cross entropy loss between the interest point detector prediction $\mathcal{X}$ on the input image, and the ground truth interest points $\mathcal{Y}$. We effectively do not apply any modifications to the interest point detection loss.
SuperPoint’s [6] descriptor loss $\mathcal{L}_{d}$, is a hinge loss between every pair of descriptors cells, $d_{h_{w}}\in\mathcal{D}$ from $\mathcal{I}$ and $d_{{h^{\prime}}{w^{\prime}}}\in\mathcal{D}^{{}^{\prime}}$ from $\mathcal{I}^{{}^{\prime}}$. "Cells" are essentially pixels in lower dimension, where for an input image of size $H\times W$, a cell has a dimension of $H_{c}=\frac{H}{8}$$\times$ $W=\frac{W}{8}$. In simpler terms, each pixel in the encoded image maps an $8\times 8$ region of the input image. SuperPoint’s descriptor loss can be defined as:
$\mathcal{L}_{d}(D,D^{\prime},S)=\frac{1}{(H_{c}W_{c})^{2}}\sum^{H_{c},W_{c}}_{\begin{subarray}{c}h=1\\ w=1\end{subarray}}\sum^{H_{c},W_{c}}_{\begin{subarray}{c}h^{\prime}=1\\ w^{\prime}=1\end{subarray}}l_{d}(d_{h_{w}},d^{\prime}_{{h^{\prime}}{w^{\prime}}};s_{hw{h^{\prime}}{w^{\prime}}})$ (7)
where $S$ is a homography induced correspondence, defined as:
$s_{hw{h^{\prime}}{w^{\prime}}}=\left\{\begin{array}[]{ll}1,&\mbox{if }||\widehat{\mathcal{H}p_{hw}}-p_{{h^{\prime}}{w^{\prime}}}||\leq 8.\\ 0,&\mbox{otherwise}\end{array}\right.$ (10)
@@
We reformulate the homography induced correspondence of SuperPoint [6], in such a way that centre pixel $p_{hw}$ can be transformed using the PrP process rather than transforming it through homography $\mathcal{H}$. Let $\mathcal{C}(\cdot)$ represent a function that takes $p_{hw}$ as an input, along with all parameters utilised in the PrP process (see Eqs. 1 and 2), for simplicity, these parameters can be combined and denoted as $M$. Therefore, the PrP induced correspondence can be defined as:
$s_{hw{h^{\prime}}{w^{\prime}}}=\left\{\begin{array}[]{ll}1,&\mbox{if }||\mathcal{C}(M;p_{hw})-p_{{h^{\prime}}{w^{\prime}}}||\leq 8.\\ 0,&\mbox{otherwise}\end{array}\right.$ (13)

### v2

@@ Abstract
Feature point detection and description is the backbone for various computer vision applications, such as Structure-from-Motion, visual SLAM, and visual place recognition. While learning-based methods have surpassed traditional handcrafted techniques, their training often relies on simplistic homography-based simulations of multi-view perspectives, limiting model generalisability. This paper presents a novel approach leveraging Neural Radiance Fields (NeRFs) to generate a diverse and realistic dataset consisting of indoor and outdoor scenes. Our proposed methodology adapts state-of-the-art feature detectors and descriptors for training on multi-view NeRF-synthesised data, with supervision achieved through perspective projective geometry. Experiments demonstrate that the proposed methodology achieves competitive or superior performance on standard benchmarks for relative pose estimation, point cloud registration, and homography estimation while requiring significantly less training data and time compared to existing approaches.
@@ 1 Introduction
Feature point detection and description under different scene viewpoints is a common starting point for many multi-view problems including Structure-from-Motion [1], visual SLAM [2, 3], or visual place recognition [4, 5]. In the past decade, different learning-based approaches to this problem [6, 7, 8, 9] have replaced handcrafted techniques in many applications [10, 11, 12]. Crucially, most of these models can be fine-tuned in a self-supervised manner on any single-view dataset. This is achieved by applying different homography warpings to the training data, simulating different viewpoints of the same scene with known point-to-point "ground-truth" mappings. While this training scheme is simple and flexible, the generated homography warpings are a crude simplification of multi-view perspectives, which can lead to limited model generalisability.
This paper aims at leveraging image synthesis with neural radiance fields (NeRFs) as a more realistic way of generating multi-view training data for feature detection and description models (see Fig. 1). Since NeRFs require multi-view data to synthesise novel views, we can no longer rely on the single image datasets typically used for training the above-mentioned homography-based methods. Therefore, we create our own dataset consisting of indoor and outdoor image sequences around static scenes and reconstruct all of them with NeRFacto [13]. This enables the generation of arbitrary viewpoints of each scene consistent with a pinhole projective model with known point-to-point mappings via point re-projection. We propose a general methodology to upgrade state-of-the-art homography-based methods to train them on projective views synthesised from NeRF-type algorithms. Our contributions are as follows:
  • We create a new multi-view dataset consisting of images from 10 different indoor and outdoor scenes, and a total of 10000 NeRF-synthesised views from these scenes with corresponding depth maps, and intrinsic and extrinsic parameters.
@@
  • We re-train adapted versions of SuperPoint [6] and SiLK [14] using our NeRF-synthesised data and compare them against the original baselines trained on the much larger MS-COCO dataset. We outperform the original baselines for relative pose estimation on ScanNet, YFCC100M and MegaDepth datasets [15, 16, 17], with similar performance for pair-wise point cloud registration, while only slightly under-performing on the HPatches homography estimation benchmark.
@@
Figure 1: Visual representation of multi-view data. Learning-based detectors and descriptors achieve supervision on single-view datasets by simulating different viewpoints through homographic warpings to the input image $I$ (see Fig. 1(a)) resulting in $I^{{}^{\prime}}$ (see Fig. 1(b)). However, we achieve supervision by directly sampling a NeRF-rendered image from a different viewpoint (see Fig. 1(c)).
@@ 2 Related Work
Research on multi-view feature detection and description has a large focus on creating representations invariant to geometric transformations (e. g. scale, rotation, affinity) and illumination conditions so that the same scene point can be reliably recognised and matched regardless of its viewpoint. Classic methods achieved this by handcrafting feature extractors that incorporate these invariance properties by design. Some of these decade-old methods stood the test of time and are still widely used today [18, 19]. More recently, the success of deep feature extraction enabled neural networks to learn these invariance properties from training data. Large open-source datasets [20, 15, 17] have been extensively employed to train either learning-based interest point detectors and descriptors [6, 21, 14, 22], or detector-free local feature matchers [23, 24, 25].
Some of these methods [6, 14] are trained in a self-supervised manner on uncalibrated RGB single-view datasets such as MS-COCO [20] by simulating multiple views as homography warpings. Other methods [21, 22] are trained with additional supervision using calibrated multi-view RGB-D datasets such as ScanNet [15]. These two classes of methods have pros and cons. The multi-view data generated by self-supervised methods is not representative of all possible viewpoint changes (homographies only model planar scenes or pure rotations). Thus it does not fully generalise to some application scenarios. On the positive side, they can be trained on an arbitrarily large number of transformations, and are also extremely flexible as they can be fine-tuned on any uncalibrated single-view dataset. Fully supervised methods are trained on more realistic multi-view data that covers full projective view changes with occlusions, however, the required training data is significantly more complex to obtain. The variety of scenes and viewpoint changes in the training data are also limited by the practicalities of real data acquisition. To get the best of both worlds, we propose using NeRF reconstructions obtained from RGB image sequences, which are both simple to acquire and enable the synthetic generation of arbitrary projective viewpoints beyond homographies.
Progress on image synthesis with NeRF and its variants [26, 27, 28] has exploded in the past years, with very fast improvements in computational efficiency, image synthesis quality, reconstruction accuracy, and input data requirements. As these techniques enable synthesising arbitrary viewpoints of a reconstructed scene, they can be used for rich data augmentation [29]. Most related to our work, they have been utilised to self-supervise multi-view methods for learning viewpoint invariant object descriptors [30], stereo disparity [31], and optical flow [32]. However, to the best of our knowledge, NeRF self-supervision has been under-explored for point feature detection and description.
@@ 3.1 NeRF Dataset
We build a NeRF dataset for 10 indoor and outdoor scenes, containing 10000 synthetic images with their corresponding intrinsic/extrinsic parameters and depth maps. We note that this dataset is extremely small compared to what is typically used to train state-of-the-art feature detection (30 times smaller than MS-COCO, 250 times smaller than ScanNet). Our goal is to show that leveraging NeRF supervision enables state-of-the-art performance on multiple benchmarks with significantly less training data.
We captured four indoor and one outdoor scenes with an iPhone 10 at 4K resolution, and one indoor and one outdoor scene with a Samsung A52 at 4k resolution. We also utilise two blender-generated indoor scenes with images sourced from [33] and one outdoor scene publicly available through [34] as they are open-source scenes that consist of high-quality RGB images. While different training sets could be constructed using existing publicly available data, we focus on a trade-off between scene diversity and minimal dataset size.
We obtain camera poses for all images using COLMAP [35] and use them to generate synthetic views. We trained a Nerfacto model for each scene, which is part of the NeRFStudio framework [13] and merges improvements from a range of NeRF implementations [36, 34, 37, 38]. For each scene, we rendered 1000 synthetic views and corresponding depth maps with $44^{\circ}$ field of view and $640\times 480$ resolution. We use NeRFStudio’s interactive real-time viewer to define custom camera trajectories for each scene that ensure diversity in viewpoints while maintaining global scene overlap. Trajectories consist of uniformly sampled smooth trajectories around the scene coupled with small rotations around the camera centre.
@@ 3.2 NeRF Point Re-Projection
$\mathbf{p^{\prime}}=\mathcal{K^{\prime}}\frac{\mathbf{P}{{}^{{}^{\prime}}}}{\mathbf{P}{{}^{{}^{\prime}}_{z}}},\quad\mathbf{P}{{}^{{}^{\prime}}}=[R^{\prime T}|-R^{\prime T}\mathbf{t^{\prime}}]\mathbf{P}.$ (2)
Assuming that reconstructed depths $d$ are accurate, the above equations enable supervising a point detector, by designing a loss function that promotes a set of re-projected points $\mathbf{p}^{\prime}$ detected in view $I$ to align with points independently detected on view $I^{\prime}$. However, many distinctive interest points lie on scene edges at a depth discontinuity. This means that even a small deviation in $\mathbf{p}$ can cause significant variations in $d$, corresponding to the depths of the background and foreground around the edges. To ensure stable re-projection around edges, we differentiate between foreground and background by focusing only on the foreground around the edges, excluding the background. To achieve this, we compute $d$ within a $5\times 5$ pixel window centred around $\mathbf{p}$. If the difference between the maximum and minimum depths within this patch does not exceed a predefined threshold $\epsilon_{d}$, we take $d$ at the interest point position. Otherwise, we use the minimum depth within the window (see example in Fig. 2).
@@ 4.1 SiLK-PrP
In the original SiLK approach, image pairs with ground-truth bi-directional correspondences are generated on the fly during training with homography warpings. We rely on our pre-rendered NeRF dataset described in section 3.1, to avoid significant computational overhead caused by on-the-fly image rendering. During training, we randomly sample image pairs $I,I^{{}^{\prime}}$ from the same scene, Eqs. 1 and 2 are then utilised during training to generate dense ground-truth point correspondences, discretised at the pixel level.
@@ 4.2 SuperPoint-PrP
SuperPoint is a self-supervised method that goes through multiple stages of training. Firstly, an encoder with a detector head is trained in a fully supervised manner to detect corners on a synthetic dataset with simple geometric shapes. The pre-trained detector, called MagicPoint, is used to generate pseudo-ground truth interest points on a real dataset (e.g. MS-COCO) using a process called homographic adaptation. This involves generating several homography warped copies of a training image, passing them through the trained detector, unwarping all predictions back to the original training image, and finally aggregating all unwarped predictions to generate the pseudo label. This process can be repeated several times, where after each round of homographic adaptation, the detector head is further trained, updating the pseudo-ground truth labels, while improving the detector head’s generalisability. After a final round of homographic adaptation, a descriptor head is added to the model, which is jointly trained with the whole model with an additional descriptor loss term on the real dataset resulting in the SuperPoint model.
With the NeRF training data approach, we use the same architecture as the original SuperPoint model. However, modify the homographic adaptation process as well as the descriptor loss to handle our NeRF training data and Point Re-Projection process.
@@ 4.2.1 Projective Adaptation
Similarly to our SiLK-PrP approach, we consider pre-rendered NeRF images as training data. In this case, we exploit the fact that training sequences for each scene are generated along a continuous trajectory.
We randomly sample from a window of 20 consecutive rendered NeRF images $I_{i\rightarrow i+19}$. We take pseudo-ground-truth points for the initial image $I_{i}$ and randomly sample a set of 14 of the remaining 19 images as the warped versions of $I_{i}$. An image from the randomly sampled set is denoted as $I_{r}$.
The probability heatmap $\mathbf{{H}_{i}}$ for image $I_{i}$ is then computed. For each image $I_{r}$ in the randomly selected set, its corresponding probability heatmap $\mathbf{{H}_{r}}$ is computed. Interest points $\mathbf{{p}_{r}}$ are then extracted from each heatmap, employing non-maximum suppression to filter out feature points that are too close to each other.
Once interest points $\mathbf{{p}_{r}}$ are obtained, the PrP process is applied to project the interest points aligning them with the initial frame ${I}_{i}$, to clarify the projected interest points from $I_{r}$ to $I_{i}$ can be denoted as $\mathbf{{p}^{{}^{\prime}}_{i}}$.
Having obtained the projected interest points $\mathbf{{p}^{{}^{\prime}}_{i}}$, we generate a mask denoted as $\mathbf{{H}_{i^{\prime}}}$, for each interest point in $\mathbf{{p}^{{}^{\prime}}_{i}}$ coordinates on the binary mask, a 3x3 patch centred around the interest point before projection is extracted from the probability heatmap $\mathbf{{H}_{r}}$, the patch is then applied on the binary heatmap $\mathbf{{H}_{i^{\prime}}}$. Thus, mask $\mathbf{{H}_{i^{\prime}}}$, serves as the probability heatmap of rendered image $I_{r}$ projected onto the input render frame ${I}_{i}$.
Finally, the input’s rendered image probability heatmap $\mathbf{{H}_{i}}$ and mask $\mathbf{{H}_{i^{\prime}}}$ for each frame, are aggregated together, non-maximum suppression is applied once more on the final aggregated heatmap to obtain pseudo-ground-truth interest points for the input rendered image $I_{i}$.
@@ 4.2.2 Descriptor Loss
SuperPoint’s [6] loss consists of an interest point detection loss $\mathcal{L}_{p}$ and a descriptor loss $\mathcal{L}_{d}$.
The interest point detection loss is a cross-entropy loss between the interest point detector prediction $\mathcal{X}$ on the input image, and the ground truth interest points $\mathcal{Y}$. We effectively do not apply any modifications to the interest point detection loss.
SuperPoint’s [6] descriptor loss $\mathcal{L}_{d}$, is a hinge loss between every pair of descriptors cells, $d_{hw}\in\mathcal{D}$ from $\mathcal{I}$ and $d_{{h^{\prime}}{w^{\prime}}}\in\mathcal{D}^{{}^{\prime}}$ from $\mathcal{I}^{{}^{\prime}}$. "Cells" are essentially pixels in the lower dimension feature map, for an input image of size $H\times W$, the feature map has a dimension of $H_{c}=\frac{H}{8}$$\times$ $W=\frac{W}{8}$. In simpler terms, each pixel in the encoded image maps an $8\times 8$ region of the input image. SuperPoint’s descriptor loss $\mathcal{L}_{d}$ can be defined as:
$\mathcal{L}_{d}(D,D^{\prime},S)=\frac{1}{(H_{c}W_{c})^{2}}\sum^{H_{c},W_{c}}_{\begin{subarray}{c}h=1\\ w=1\end{subarray}}\sum^{H_{c},W_{c}}_{\begin{subarray}{c}h^{\prime}=1\\ w^{\prime}=1\end{subarray}}l_{d}(d_{hw},d^{\prime}_{{h^{\prime}}{w^{\prime}}};s_{hw{h^{\prime}}{w^{\prime}}})$ (3)
where $l_{d}$ is defined as:
$l_{d}(d,d^{\prime};s)=\lambda_{d}*s*max(0,m_{p}-d^{T}d^{\prime})+(1-s)*max(0,d^{T}d^{\prime}-m_{n})$ (4)
where $m_{p}$ is the hinge loss’s positive margin, $m_{n}$ is the negative margin and $\lambda_{d}$ balances between the positive and negative correspondences. Additionally, $S$ in Eq. 3 is a homography-induced correspondence, defined as:
$s_{hw{h^{\prime}}{w^{\prime}}}=\left\{\begin{array}[]{ll}1,&\mbox{if }||\widehat{\mathcal{H}p_{hw}}-p_{{h^{\prime}}{w^{\prime}}}||\leq\epsilon_{s}.\\ 0,&\mbox{otherwise}\end{array}\right.$ (7)
@@
We reformulate the homography-induced correspondence of SuperPoint [6], in such a way that centre pixel $p_{hw}$ can be transformed using the PrP process rather than transforming it through homography $\mathcal{H}$. Let $\mathcal{C}(\cdot)$ represent a function that takes $p_{hw}$ as an input, along with all parameters utilised in the PrP process (see Eqs. 1 and 2), for simplicity, these parameters can be combined and denoted as $M$. Therefore, the PrP-induced correspondence can be defined as:
$s_{hw{h^{\prime}}{w^{\prime}}}=\left\{\begin{array}[]{ll}1,&\mbox{if }||\mathcal{C}(M;p_{hw})-p_{{h^{\prime}}{w^{\prime}}}||\leq\epsilon_{s}.\\ 0,&\mbox{otherwise}\end{array}\right.$ (10)